OpenStudy (anonymous):

solve using elimination method: 2x – y = 8, y = 5x + 3

OpenStudy (anonymous):

do I change 2nd equation to : 5x -y =3? I am confused

jhonyy9 (jhonyy9):

so yes you are right and after this for we can eliminet the term of y so for we need multiply first equation by -1 and after this you assum thses two equations

jhonyy9 (jhonyy9):

2x-y=8 5x-y=-3

jhonyy9 (jhonyy9):

sorry 5x-y=-3 will be correct

OpenStudy (anonymous):

so, how do I eliminate for the 'y' variable? do I multiply the reciprocal of one 'y' variable to one equation? Like for the 1st equation I would multiply '1' to eliminate the 'y'?

OpenStudy (anonymous):

oh wait...I meant '-1' instead of 1

jhonyy9 (jhonyy9):

so 2x-y=8 5x-y=-3 multiply by -1 2x-y=8 -5x+y=3 ----------- -3x 0 =11 3x= - 11 x= -11/3

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

and now calcule the y

OpenStudy (anonymous):

then I get -2x + y = -8, then I add or subtract to 5x - y = -3? ---oh ok tyty

jhonyy9 (jhonyy9):

and now you can calcule the y yes ?

jhonyy9 (jhonyy9):

camon i will wait it

jhonyy9 (jhonyy9):

so if right ok than thank you the medal good luck bye

jhonyy9 (jhonyy9):

so if right ok than thank you the medal good luck bye

jhonyy9 (jhonyy9):

so y= 5x+3 y=5(-11/3) +3 =-55/3 +3 = (-55+9)/3 = -46/3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yes thank you!

jhonyy9 (jhonyy9):

ok medal please

OpenStudy (anonymous):

are you helping out people just to get a medal? that's a little wrong. you should help others out in the spirit of helpfulness, not greed!

jhonyy9 (jhonyy9):

no just for medal but just than the exersice will be signed solved right sure

jhonyy9 (jhonyy9):

or you need again learning on this website how make the things